OpenStudy (anonymous):

Use the distributive property to simplify the expression below -5x+15 over -5 -x - 3 x + 15 x + 3 x - 3

OpenStudy (anonymous):

(-5x + 15)/ -5= -5( x - 3) / -5 = x-3 so option D